Use the method of Lagrange Multipliers to find the points on the surface ?? − ? 2 = 9 that are closest to the origin
Answer #1

Ando Р (wy, ) Pluz, 2) Р O (دردره ) so Let gren surface xy-2², and the have to find bont on the surface of that curre kinosemultibly (ty and (ii): 4 xy = x xy = xy(4-x²)=0. so eitners ayzo or 4- x 20 = 7212, 4 If xy =o, then but into (ii): ny-2² = 9I but x=-4 in ay-z²9 =) - 7²0² = 9 quel Z (roreal value of of y) but h=4 and 220 in ny-z = 9 = 4²²9 =) y=+3. < = 3 anel z= 0.